OpenStudy (anonymous):

solve by completing the square: 3x^2-9x-12=0

OpenStudy (anonymous):

\[3x^2-9x-12=0\] divide by 3 get \[x^2-3x-4=0\] add 4 get \[x^2-3x=4\] take half of three get \[(x-\frac{3}{2})^2=4+\frac{9}{4}=\frac{25}{4}\] then take the square root of both sides get \[x-\frac{3}{2}=\pm\frac{5}{2}\] and finally \[x=\frac{3}{2}+\frac{5}{2}=\frac{8}{2}=4\] pr \[x=\frac{3}{2}-\frac{5}{2}=\frac{-2}{2}=-1\]

OpenStudy (anonymous):

it would have been tonnes easier to factor as \[(x+1)(x-4)=0\] etc

OpenStudy (anonymous):

youre the best!!

OpenStudy (anonymous):

ty (blush)